# Compositional Abilities Emerge Multiplicatively: Exploring Diffusion Models on a Synthetic Task

### v4

@@ 1 Introduction
The scaling of data, models, and computation has unleashed powerful capabilities in generative models, enabling controllable synthesis of realistic images [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11], 3D scenes [12, 13, 14, 15, 16], videos [17, 18, 19, 20, 21, 22, 23, 24, 25], accurate image-editing [26, 27, 28, 29, 30, 31], and semantically coherent text generation [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. With increased interest in incorporating these models in day-to-day applications [47, 48, 49, 50], e.g., to improve robotic systems via better planning and grounding [51, 52, 53, 54, 55, 56, 57, 58, 59], analyzing and improving their reliability has become crucial.
@@ 3 Concept Graph: A Minimalistic Framework for Compositionality
[…]
@@
[…]
@@
[…]
@@ 3.1 Experimental and Evaluation Setup
[…]
@@ 4 Multiplicative Emergence of Compositional Abilities
Learning dynamics respect the structure of the concept graph (Fig. 5). We find the ability to compositionally generalize, i.e., produce samples from out-of-distribution concept classes, emerges at a rate which is inversely related to a class’s concept distance with respect to classes seen in training. Specifically, in Fig. 5 (a) we show the learning dynamics of the model, where lightblue nodes denote concept classes within the training dataset, while pink and darkpink nodes respectively denote classes at concept distances of 1 and 2 from classes in the training dataset. As the model learns to fit its training data (lightblue nodes), it learns capabilities that can be composed to produce samples from out-of-distribution concept classes (pink / darkpink nodes). Fig. 5 (b) further shows that the learning dynamics of compositional generalization are influenced by the concept distance from the training set: the model first learns the concept classes in the training dataset (lightblue lines) and then generalizes to concept classes with a concept distance of 1 from the training dataset (pink lines). Thereafter, the model suddenly acquires the capability to compositionally generalize to a concept class with a concept distance of 2 from the training dataset (darkpink line). Fig. 5 (c) shows the images generated by the model over time. We observe that rough shapes and sizes are learned relatively early in training, by the 4th epoch, while the color is highly biased to be red, the majority color in the training dataset, up to the 10th epoch. Then, around the 20th epoch, the model learns to generate the underrepresented
@@
[…]
Multiplicative influence of capabilities drives the sudden emergence of compositional generalization (Fig. 6). The interpretability of our experimental setup illustrates that the multiplicative reliance on underlying capabilities is the critical structure behind the sudden emergence of compositional generalization in our setup. For example, in Fig. 6 (a) we show the dynamics of an additive vs. multiplicative accuracy measure for generating concept class {111, (triangle, small, blue)}, which has a concept distance of 2 from the training data. We observe a sudden improvement of the multiplicative measure. To better understand the above, in Fig. 6 (b), we plot the accuracy of probes used for predicting each concept variable (shape, size, color). From the plot, we notice that the model fails to acquire strong color transformation capabilities until the final stage of training, effectively bottlenecking compositional generalization of $111$ class. This empirical observation leads us to the following hypothesis on emergence of compositional abilities in generative models.
@@
Toy analysis. We consider a toy compositional problem which demonstrates how complex capabilities can rapidly develop (“emerge”) from a set of atomic abilities. Assume there are $n$ atomic abilities, each with a probability $p$ of being learned in a given time step, i.e., the dynamics of learning an ability can be modeled as a Bernoulli coin flip: once the coin turns from 00 to $1$, the model learns the said ability. The probability that the ability will be learned in $t$ steps is thus given by $1-(1-p)^{t}$. An implicit assumption in this model is that the learning dynamics of atomic abilities are independent (hence the name atomic) and that once an atomic ability is learned, it will not be forgotten. Now, our goal is to characterize the dynamics of learning a capability that involves composing these atomic abilities. The multiplicative emergence hypothesis argues that the learning such a capability by time $t$ requires that the model have learned the $n$ atomic abilities by that time. We then get the probability that the compositional capability has been learned by time $t$ is $P(n)=\left(1-(1-p)^{t}\right)^{n}$. These dynamics are plotted in Fig. 7 (a) and look strikingly similar to the rapid learning (emergence) we found in our diffusion model setup (see Fig. 5 b)): for increasing degree of compositionality, we see much sharper learning dynamics (i.e., emergence) that is controlled by the concept distance w.r.t. atomic abilities. Further, assuming a given threshold probability $P^{*}$ at which one claims a capability has been learned, for a degree of composition $n$, we can solve for the critical time $t^{*}$ at which a compositional capability is learned as follows: $t^{*}=\left\lceil\frac{\log(1-(P^{*})^{1/n})}{\log(1-p)}\right\rceil$. Plotting this in Fig. 7 (b), we see the progress on increasingly more compositional tasks is logarithmic w.r.t. the number of atomic concepts being composed (aka concept distance). Importantly, this implies if we allow the model to train infinitely and it learns several atomic abilities, it will have an explosion of capabilities due to the inherent compositionality of the data generating process—we further investigate this in a follow up work [118].
@@ 4.1 Challenges for Compositional Generalization
Figure 8: How does the frequency of data samples impact the learning of capabilities? We systematically control the frequency of a specific concept class in the training dataset and observe how it affects the model’s learning of capabilities. (a) Capability to alter colors is quickly learned after introducing approx. 10 samples with a concept class of large blue circle, $001$. (b) In contrast, a critical threshold (marked with dotted vertical lines) exists for learning a capability to alter the shape: as we gradually introduce samples, with a concept class of large red triangle, $100$.
Critical frequency for learning capabilities (Fig. 8). We first probe the effect of changing the frequencies of samples from different concept classes in the training data and examine how this affects the model’s ability to learn and compose capabilities involving that concept class. Results are shown in Fig. 8 and demonstrate how the frequency of color and size concepts in the training data impacts the generalization capabilities of the diffusion model. Specifically, we change the number of samples in the training data from 0 to 300 for concept class $001$ (Fig. 8 (a)) and class $100$ (Fig. 8 (b)). As can be seen, low frequencies of certain concept values degrade the accuracy of the model in both settings. Notably, training for out-of-distribution concept classes (pink lines) requires more samples than that for in-distribution ones (lightblue lines). This suggests that as the sample size grows, memorization occurs first, and generalization is achieved superlinearly as a function of data frequency. More importantly, we observe a critical number of samples are required before we can see the onset of capabilities to alter a concept. Specifically, in Fig. 8 (a), we can see that the model rapidly learns the color concept after being provided with 10 samples for a concept of large blue circle, $001$. In contrast, in Fig. 8 (b), the model learns the shape concept only after reaching a certain threshold in the number of samples with a concept of large red triangle, $100$.
@@
Figure 9: Fine-tuning is not a remedy for misgeneralization in an adversarial task. (a) The model struggles to generalize when the training setup is configured adversarially. As shown in the plot, the model incorrectly produces triangles (instead of circles) for both $011$ and $010$. (b) The model faces difficulty learning new concepts through fine-tuning. We add node $101$ (triangle, large, blue) to the dataset and attempt to fine-tune the model. However, even with a large learning rate equal to the one used for training, the model fails to learn the capability to alter object size.
Fine-tuning is not a remedy for misgeneralization in an adversarial task. (Fig. 9). We next evaluate a setting where concept classes present in the training data are not neighbors, i.e., their concept distances are greater than 1, but nonetheless represent all concept variables to allow a model to learn relevant capabilities. Specifically, as shown in Fig. 9 (a), we use only the following four pairs of tuples and images for training: $\{000,(\text{circle},\text{red},\text{large})\}$, $\{100,(\text{triangle},\text{red},\text{large})\}$, $\{001,(\text{circle},\text{blue},\text{small})\}$, $\{111,(\text{triangle},\text{blue},\text{small})\}$. Arguably, capabilities corresponding to shape change ($000$ to $100$) and color change ($000$ to $001$) can still be learned as the setup for these is similar to our prior experiments. However, our results show this yields an interesting failure mode; specifically, in Fig. 9 (a) we find the model fails to dissociate a second element value of 1 with the shape of a small triangle, and this bias causes the model to generate small triangles for both $011$ and $010$, when it should have produced small circles.
We further analyze if the failure above can be fixed via mere downstream fine-tuning (see Fig. 9 (b)). Specifically, we fine-tune the trained model on a dataset that includes the concept class of $101$ (left), $010$ (middle), and $011$ (right). We find the model still generates small triangles for $011$ and $010$ even after fine-tuning. When the concept classes $010$ (middle) and $011$ (right) are added, the newly introduced concepts “overwrite” all existing concepts, causing the previously learned concepts (e.g., the color and shape concept for $101$) to be forgotten. These results present a further challenge in addressing learned bias.
@@ 5 Conclusion
Authors’ Contributions
ESL initiated the research by defining the problem. ESL and HT formulated the theoretical framework, the concept graph, and collaborated with MO on the experimental design. MO conducted most experiments and, along with ESL and HT, produced the visualizations and figures. ESL and HT authored the introduction. ESL was responsible for the literature review and related work. MO, ESL, and HT jointly worked on the results and discussion sections. Throughout the study, MO, ESL, and HT collaborated on analyzing experimental results and refining the manuscript. HT and RPD provided supervisory guidance and feedback.
@@ A.6 Experiment with real data
We sourced our data from the CelebA dataset [^3] https://mmlab.ie.cuhk.edu.hk/projects/CelebA.html. Our training dataset consists of 15,000 facial images, each labeled with its respective concept class. We delineated the concept classes based on these attributes: gender (classified as "Male" or otherwise), facial expression ("Smiling" or not), and hair color (either "Black Hair" or "Brown Hair"). We adjusted the images to a resolution of $48\times 48$ pixels.
@@ Appendix B Additional experimental results
[…]
In Fig. 17, we investigate the bottlenecks that affect compositional generalization by examining the prediction accuracy of concept variables such as shape, color, and size, as well as the overall classification accuracy. To evaluate this, we generated 50 samples in each epoch of training and classified them using a linear probe. Our findings reveal that various concept bottlenecks impede the learning process for different compositional objects. For instance, the generalization of blue-colored objects ($011$, $101$, $111$) is hindered by the color concept, indicating that the model struggles to correctly identify objects based on their color when it comes to these particular compositions. Similarly, the generalization to small red triangles ($111$) is bottlenecked by the shape concept, implying that the model faces difficulties in accurately recognizing the shape of these specific compositions. These observations highlight the challenges associated with compositional generalization and shed light on the specific concept bottlenecks that hinder the learning process. Understanding and addressing these bottlenecks can contribute to the development of more robust models capable of achieving better compositional generalization in various contexts.
Figure 17: Analyzing the bottlenecks of compositional generalization. We examine the prediction accuracy of concept variables (shape, color, and size) and the overall classification accuracy. In each epoch of training, 50 samples were generated and classified using linear probe. We observe that various concept bottlenecks hinder the learning of different compositional objects. For example, compositional generalization to blue-colored objects ($011$, $101$, $111$) is bottlenecked by the color concept, while generalization to small red triangles ($111$) is bottlenecked by the shape concept.
@@
In Fig. 21, we conducted experiments with different configurations by using varied sets of nodes in the training data. In Fig. 21(a) and Fig. 21(b), we inverted the axes of the concept graph from Fig. 9(a): (a) from bottom to top and (b) from left to right. In Fig. 21(c) and Fig. 21(d), we inverted the axes of the concept graph from Fig. 5(a): (c) from bottom to top and (d) from left to right. In Fig. 21(b) and Fig. 21(c), we observe that the model incorrectly generates triangles for both $011$ and $010$, even though circles are expected. In the other cases, the model outputs correct images for all the nodes. Our observations indicate that the model struggles to to dissociate a second element value of 1 with the shape of a small triangle, exhibiting a bias to generate triangles rather than circles for small objects.

### v5

@@ 1 Introduction
The scaling of data, models, and computation has unleashed powerful capabilities in generative models, enabling controllable synthesis of realistic images [1, 2, 3, 4, 5, 6, 7, 8, /, 9, 10], 3D scenes [11, 12, 13, 14, 15], videos [16, 17, 18, 19, 20, 21, 22, 23, 24], accurate image-editing [25, 26, 27, 28, 29, 30], and semantically coherent text generation [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. With increased interest in incorporating these models in day-to-day applications [46, 47, 48, 49], e.g., to improve robotic systems via better planning and grounding [50, 51, 52, 53, 54, 55, 56, 57, 58], analyzing and improving their reliability has become crucial.
@@ 3 Concept Graph: A Minimalistic Framework for Compositionality
Figure 3: Concept graphs. We organize our study using a simple but expressive framework called concept graphs. The basis of a concept graph is a set of primitives called concept variables (e.g., $\mathtt{shape}\text{ }$, $\mathtt{color}\text{ }$, etc.). A subset of these variables is instantiated with specific values to yield a concept class, e.g., $\{\mathtt{shape=0,size=0,color=1}\}$ implies a small, blue circle. This is akin to defining a broad set of objects that share some common properties, e.g., all lizards in Fig. 1 belong to the same species, despite their diverse colors. A specific object in this class is instantiated by fixing the remaining variables, e.g., small, blue circles at different locations. Each concept class corresponds to a graph node, where nodes are connected if their concept classes differ by a concept distance of 111.
@@
For instance, for geometric objects shown in Fig. 3, concept variables may include $\mathtt{shape}\text{ }$, $\mathtt{color}\text{ }$, $\mathtt{size}\text{ }$, $\mathtt{location}\text{ }$, and $\mathtt{angle}$. These variables take on values from a pre-specified range, called concept values, as described below.
@@
Note that the remaining $n-p$ concept variables are free in the above definition. When these free variables are specified, we get a specific object from our data-generating process. That is, a concept class represents a family of objects by specifying the values of a pre-defined subset of concept variables. For example, in Fig. 1, different colored lizards instantiate images (objects) from the species (concept class) of lizards; here, color of the lizard serves as a free variable. Similarly, in the geometric objects scenario of Fig. 3, a “small red circle” would be a concept class for which the $\mathtt{shape}\text{ }$, $\mathtt{color}\text{ }$, and $\mathtt{size}\text{ }$variables have been assigned specific values. Objects are images designated by associating precise values with the remaining concept variables of $\mathtt{location}\text{ }$and $\mathtt{angle}\text{ }$. Next, we introduce the notion of concept distance, which serves as a proxy to succinctly describe the dissimilarity between two concept classes.
@@ 3.1 Experimental and Evaluation Setup
Evaluation Metric. Evaluating whether a generated image corresponds to the desired concept class might require a human in the loop. To eliminate this requirement, we follow methods from literature on disentanglement and train classifier probes to predict whether a generated image possesses some property of interest [103, 109, 110, 111, 112, 113, 114]. Specifically, we use the diffusion model training data to train three linear probes, which respectively infer the following three concept variables: $\mathtt{shape}\text{ }$, $\mathtt{color}\text{ }$, and $\mathtt{size}\text{ }$. We define a model’s accuracy for generating images of a given concept class as the product of the probabilities outputted by the three probes that each concept variable matches the value for the concept class. Note that a random classifier for a concept variable will have an accuracy of 0.5. We indicate this random baseline with dotted, gray lines in our plots when useful.
@@ 4 Multiplicative Emergence of Compositional Abilities
Learning dynamics respect the structure of the concept graph (Fig. 5). We find the ability to compositionally generalize, i.e., produce samples from out-of-distribution concept classes, emerges at a rate which is inversely related to a class’s concept distance with respect to classes seen in training. Specifically, in Fig. 5 (a) we show the learning dynamics of the model, where lightblue nodes denote concept classes within the training dataset, while pink and darkpink nodes respectively denote classes at concept distances of 1 and 2 from classes in the training dataset. As the model learns to fit its training data (lightblue nodes), it learns capabilities that can be composed to produce samples from out-of-distribution concept classes (pink / darkpink nodes). Fig. 5 (b) further shows that the learning dynamics of compositional generalization are influenced by the concept distance from the training set: the model first learns the concept classes in the training dataset (lightblue lines) and then generalizes to concept classes with a concept distance of 1 from the training dataset (pink lines). Thereafter, the model suddenly acquires the capability to compositionally generalize to a concept class with a concept distance of 2 from the training dataset (darkpink line). Fig. 5 (c) shows the images generated by the model over time. We observe that rough shapes and sizes are learned relatively early in training, by the 4th epoch, while the color is highly biased to be red, the majority color in the training dataset, up to the 10th epoch. Then, around the 20th epoch, the model learns to generate the underrepresented.
@@
color (blue) for concept classes with a concept distance of 1. Finally, around the 40th epoch, the model learns to generate the underrepresented color (blue) for the class at distance 2, showing a sudden emergence of capability to generate samples from that class. The observations above generalize to higher-dimensional concept graphs as well (see Fig. 16). Specifically, we now introduce a fourth concept variable $\mathtt{backgroundcolor}\text{ }$={white, black} and again find that compositional generalization occurs in the order governed by the distance from the training set: the accuracy begins to rise for concept distance of 1, continues to increase for distance 2, and peaks for 3.
Multiplicative influence of capabilities drives the sudden emergence of compositional generalization (Fig. 6). The interpretability of our experimental setup illustrates that the multiplicative reliance on underlying capabilities is the critical structure behind the sudden emergence of compositional generalization in our setup. For example, in Fig. 6 (a) we show the dynamics of an additive vs. multiplicative accuracy measure for generating concept class {111, (triangle, small, blue)}, which has a concept distance of 2 from the training data. We observe a sudden improvement of the multiplicative measure. To better understand the above, in Fig. 6 (b), we plot the accuracy of probes used for predicting each concept variable (shape, size, color). From the plot, we notice that the model fails to acquire strong color transformation capabilities until the final stage of training, effectively bottlenecking compositional generalization of 111111111 class. This empirical observation leads us to the following hypothesis on emergence of compositional abilities in generative models.
@@
Toy analysis. We consider a toy compositional problem which demonstrates how complex capabilities can rapidly develop (“emerge”) from a set of atomic abilities. Assume there are $n$ atomic abilities, each with a probability $p$ of being learned in a given time step, i.e., the dynamics of learning an ability can be modeled as a Bernoulli coin flip: once the coin turns from 0 to 111, the model learns the said ability. The probability that the ability will be learned in $t$ steps is thus given by $1-(1-p)^{t}$. An implicit assumption in this model is that the learning dynamics of atomic abilities are independent (hence the name atomic) and that once an atomic ability is learned, it will not be forgotten. Now, our goal is to characterize the dynamics of learning a capability that involves composing these atomic abilities. The multiplicative emergence hypothesis argues that the learning such a capability by time $t$ requires that the model have learned the $n$ atomic abilities by that time. We then get the probability that the compositional capability has been learned by time $t$ is $P(n)=\left(1-(1-p)^{t}\right)^{n}$. These dynamics are plotted in Fig. 7 (a) and look strikingly similar to the rapid learning (emergence) we found in our diffusion model setup (see Fig. 5 b)): for increasing degree of compositionality, we see much sharper learning dynamics (i.e., emergence) that is controlled by the concept distance w.r.t. atomic abilities. Further, assuming a given threshold probability $P^{*}$ at which one claims a capability has been learned, for a degree of composition $n$, we can solve for the critical time $t^{*}$ at which a compositional capability is learned as follows: $t^{*}=\left\lceil\frac{\log(1-(P^{*})^{1/n})}{\log(1-p)}\right\rceil$. Plotting this in Fig. 7 (b), we see the progress on increasingly more compositional tasks is logarithmic w.r.t. the number of atomic concepts being composed (aka concept distance). Importantly, this implies if we allow the model to train infinitely and it learns several atomic abilities, it will have an explosion of capabilities due to the inherent compositionality of the data generating process—we further investigate this in a follow up work [118].
@@ 4.1 Challenges for Compositional Generalization
Figure 8: How does the frequency of data samples impact the learning of capabilities? We systematically control the frequency of a specific concept class in the training dataset and observe how it affects the model’s learning of capabilities. (a) Capability to alter colors is quickly learned after introducing approx. 10 samples with a concept class of large blue circle, 001001001. (b) In contrast, a critical threshold (marked with dotted vertical lines) exists for learning a capability to alter the shape: as we gradually introduce samples, with a concept class of large red triangle, 100100100.
Critical frequency for learning capabilities (Fig. 8). We first probe the effect of changing the frequencies of samples from different concept classes in the training data and examine how this affects the model’s ability to learn and compose capabilities involving that concept class. Results are shown in Fig. 8 and demonstrate how the frequency of color and size concepts in the training data impacts the generalization capabilities of the diffusion model. Specifically, we change the number of samples in the training data from 0 to 300 for concept class 001001001 (Fig. 8 (a)) and class 100100100 (Fig. 8 (b)). As can be seen, low frequencies of certain concept values degrade the accuracy of the model in both settings. Notably, training for out-of-distribution concept classes (pink lines) requires more samples than that for in-distribution ones (lightblue lines). This suggests that as the sample size grows, memorization occurs first, and generalization is achieved superlinearly as a function of data frequency. More importantly, we observe a critical number of samples are required before we can see the onset of capabilities to alter a concept. Specifically, in Fig. 8 (a), we can see that the model rapidly learns the color concept after being provided with 10 samples for a concept of large blue circle, 001001001. In contrast, in Fig. 8 (b), the model learns the shape concept only after reaching a certain threshold in the number of samples with a concept of large red triangle, 100100100.
@@
Figure 9: Fine-tuning is not a remedy for misgeneralization in an adversarial task. (a) The model struggles to generalize when the training setup is configured adversarially. As shown in the plot, the model incorrectly produces triangles (instead of circles) for both 011011011 and 010010010. (b) The model faces difficulty learning new concepts through fine-tuning. We add node 101101101 (triangle, large, blue) to the dataset and attempt to fine-tune the model. However, even with a large learning rate equal to the one used for training, the model fails to learn the capability to alter object size.
[…]
We further analyze if the failure above can be fixed via mere downstream fine-tuning (see Fig. 9 (b)). Specifically, we fine-tune the trained model on a dataset that includes the concept class of 101101101 (left), 010010010 (middle), and 011011011 (right). We find the model still generates small triangles for 011011011 and 010010010 even after fine-tuning. When the concept classes 010010010 (middle) and 011011011 (right) are added, the newly introduced concepts “overwrite” all existing concepts, causing the previously learned concepts (e.g., the color and shape concept for 101101101) to be forgotten. These results present a further challenge in addressing learned bias.
@@ A.6 Experiment with real data
We sourced our data from the CelebA dataset [^3] https://mmlab.ie.cuhk.edu.hk/projects/CelebA.html. Our training dataset consists of 15,000 facial images, each labeled with its respective concept class. We delineated the concept classes based on these attributes: gender (classified as "Male" or otherwise), facial expression ("Smiling" or not), and hair color (either "Black Hair" or "Brown Hair"). We adjusted the images to a resolution of $\times 48$ 48 pixels.
@@ Appendix B Additional experimental results
In Fig. 16, we generalize the findings in Fig. 5 to a larger concept graph, i.e., one with four concept variables. Specifically, we now introduce a fourth concept variable $\mathtt{backgroundcolor}\text{ }$={white, black} and again find that compositional generalization occurs in the order governed by the distance from the training set: the accuracy begins to rise at a concept distance of 1, continues to increase at a distance of 2, and peaks at 3.
In Fig. 17, we investigate the bottlenecks that affect compositional generalization by examining the prediction accuracy of concept variables such as shape, color, and size, as well as the overall classification accuracy. To evaluate this, we generated 50 samples in each epoch of training and classified them using a linear probe. Our findings reveal that various concept bottlenecks impede the learning process for different compositional objects. For instance, the generalization of blue-colored objects (011011011, 101101101, 111111111) is hindered by the color concept, indicating that the model struggles to correctly identify objects based on their color when it comes to these particular compositions. Similarly, the generalization to small red triangles (111111111) is bottlenecked by the shape concept, implying that the model faces difficulties in accurately recognizing the shape of these specific compositions. These observations highlight the challenges associated with compositional generalization and shed light on the specific concept bottlenecks that hinder the learning process. Understanding and addressing these bottlenecks can contribute to the development of more robust models capable of achieving better compositional generalization in various contexts.
Figure 17: Analyzing the bottlenecks of compositional generalization. We examine the prediction accuracy of concept variables (shape, color, and size) and the overall classification accuracy. In each epoch of training, 50 samples were generated and classified using linear probe. We observe that various concept bottlenecks hinder the learning of different compositional objects. For example, compositional generalization to blue-colored objects (011011011, 101101101, 111111111) is bottlenecked by the color concept, while generalization to small red triangles (111111111) is bottlenecked by the shape concept.
@@
In Fig. 21, we conducted experiments with different configurations by using varied sets of nodes in the training data. In Fig. 21(a) and Fig. 21(b), we inverted the axes of the concept graph from Fig. 9(a): (a) from bottom to top and (b) from left to right. In Fig. 21(c) and Fig. 21(d), we inverted the axes of the concept graph from Fig. 5(a): (c) from bottom to top and (d) from left to right. In Fig. 21(b) and Fig. 21(c), we observe that the model incorrectly generates triangles for both 011011011 and 010010010, even though circles are expected. In the other cases, the model outputs correct images for all the nodes. Our observations indicate that the model struggles to to dissociate a second element value of 1 with the shape of a small triangle, exhibiting a bias to generate triangles rather than circles for small objects.